two integers have a difference of -11, and a sum of -3What are they?
tickets for a benefit went on sale for three weeks, twuice as many sold in week three than first two weeks combined. A total of 1095 sold. How many sold the thirf week?

Answer by rfer(12657)

1095/3=365*2=730